### v2

@@ 4.5 Summary
This section measures the encryption practices of consumer IoT devices. The study reveals that Chinese manufacturers adopt safe encryption protocols for device communications, and encrypted traffic dominates most communications. However, some devices still use deprecated protocols, though their traffic constitutes a small proportion(7.6%), and some devices use proprietary protocols developed by the manufacturers rather than standardized protocols. bringing about certain security risks. Encouragingly, following firmware updates, 10 devices adopted more secure encryption protocols after the update, indicating a positive trend toward encryption protocol usage. Most devices use secure certificates, but prevalent self-signed certificates and insecure algorithms persist as potential risks. In terms of certificate verification, many devices have the problem of insecure certificate verification mechanisms, which may also bring potential risks. While direct PII exposure in unencrypted traffic is rare, inferred user preferences could enable targeted advertising. These findings indicate that the devices should adopt standard encryption protocols and reduce reliance on self-signed certificates to further minimize risks.
@@ 5 Comparisons on Different Regions
This section compares consumer IoT device traffic measurement results in China with those in other regions. We choose Ren et al.[5] and Paracha M. T. et al.[13] for comparison, as their studies align with our research purpose, focusing on traffic destinations and encryption analysis.
@@ 5.1 Traffic Destinations
We have summarized the differences of traffic destinations between China and other regions in the table 6. In terms of traffic destinations, compared to consumer IoT devices in the United States and the United Kingdom, Chinese devices predominantly communicate domestically.
@@
Other Reigions 0.10% 0.10% 3.38% Greatbit Microsoft Microsoft
@@
party China (77) United States (46) United Kingdom( 35) Tencent Kingsoft Kingsoft
@@
    The ratio between the United States and the United Kingdom was estimated in Ren et al.’s work.
According to Ren et al. [5], the 35 devices in the United Kingdom communicated with 7 overseas regions. Approximately 70% of the United Kingdom’s device traffic flowed outside their regions (primarily to the United States). In contrast, although Chinese consumer IoT devices contacted 18 overseas countries, overseas traffic only accounted for 3% of total bytes. This highlights Chinese consumer IoT devices’ stronger reliance on domestic services than their United Kingdom counterparts. The devices in the United Kingdom partially rely on overseas cloud platforms and service providers (such as AWS, Google Cloud, and Akamai), resulting in frequent cross-border transmission of their data to overseas data centres. This increases the risks of data leakage. In contrast, most data transmission activities of Chinese consumer IoT devices occur within the country, indicating a high degree of local service deployment. This phenomenon reflects the high degree of reliance of Chinese manufacturers on local services.
Regarding commonly contacted organizations, Chinese consumer IoT devices predominantly interact with domestic providers. In the Ren et al. study [5], the United Kingdom devices are mainly connected to the United States organizations, such as Amazon, Google, and Akamai. In our analysis, except for Google, Chinese devices primarily communicated with domestic organizations such as Alibaba Cloud, Xiaomi, and Tencent. These providers offer diverse services, including IoT platform services, cloud storage services, and video streaming and media content (See Section 3.2 for details).
@@ 5.2 Encryption Analysis
Encryption practices vary regionally, but unencrypted traffic proportions remain low overall. Ren et al. [5] proposed entropy-based encryption detection, reporting unencrypted traffic proportions below 13% for devices in the United States and the United Kingdom. Using similar methods, we come to a similar conclusion: the percentage of unencrypted traffic of consumer IoT devices in China is low, with a peak of 10%.
Regarding encryption protocol usage, TLS 1.2 remains the most widely adopted, while TLS 1.3 adoption is still limited. Paracha M. T. et al. [13] analyzed 40 consumer IoT devices in the United States in 2021 and found that TLS 1.2 is the dominant protocol, with minimal use of TLS 1.3 and deprecated versions. Our findings align with these results. Among 70 Chinese devices, 82% utilized TLS 1.2, while 7% still relied on insecure protocols (TLS 1.1/SSL), and 14% employed proprietary protocols developed by manufacturers. For certificate validation, Paracha M. T. et al. [13] examined 32 devices and found that 11 were vulnerable to certificate replacement. In our study, 17 out of 40 devices generated error messages when the certificate was replaced, while 23 devices exhibited no errors or anomalies, indicating that certificate verification remains a security concern.
In terms of PII exposure, Ren et al.’s study [5] identified PII in unencrypted traffic, including MAC addresses, UUIDs, device IDs, geolocation data, and user-assigned device names. In contrast, Chinese devices exposed minimal PII—primarily singer names in speaker’s traffic, indirectly revealing music preferences. This indicates stronger privacy protection in Chinese devices’ unencrypted traffic.
@@ 5.3 Summary
A clear regional difference is observed in the distribution of traffic destinations and encryption practices. Chinese consumer IoT devices are more reliant on local service providers, in contrast to devices in the United Kingdom, which predominantly rely on international cloud platforms such as Amazon. This suggests that China has developed a more self-reliant and locally integrated IoT ecosystem, with limited dependence on international platforms. Encryption practices and certificate implementations are generally consistent across regions, ensuring relative security, but a certain proportion of devices still use outdated protocols across regions, which could pose potential security and privacy risks. Meanwhile, there are weaknesses with certificate verification in devices from different regions, posing a security concern. The PII contained in the unencrypted traffic of Chinese consumer IoT devices is less than that of their counterparts in the United Kingdom and the United States.
@@ 6 Related Work
The Measurement of IoT traffic. Researchers have conducted a lot of measurements on IoT traffic, mainly focusing on security and privacy. Security measurements involve mapping IoT backend servers [17], analyzing the impact of botnets on devices [26, 27], and investigating data leakage and device vulnerabilities through companion apps [29, 31, 32, 33]. Said et al. [17] mapped IoT backend servers by analyzing ISP passive traffic data, revealing relationships between backend providers. Arman et al. [26] assessed the impact of two ISP security policies on the Mirai botnet. Stephen et al. [27] studied the Hajime botnet’s effect on devices using active scanning and passive DNS backscatter traffic collection. These studies primarily focus on device-level evaluations while overlooking the user perspective. Privacy measurements include traffic destination identification [30], user data collection analysis [23, 24], and the effectiveness of existing privacy protection policies [25]. Mandalari et al. [30] proposed an automated method to identify IoT traffic destinations, distinguishing between essential and non-essential ones. Aniketh et al. [23] examined privacy leaks in local communication models using standard protocols. Dubois et al. [24] investigated how, when, and why smart speakers record external audio, while Anna Maria et al. [25] tested the effectiveness of existing privacy protection measures. Compared with these studies, our work directly analyzes IoT device security practices by evaluating traffic destinations and encryption practices. Ren et al. [5] and Paracha et al. [13] have similar research purposes with ours. Muhammad Talha Paracha et al. [13] evaluated the security of TLS connections in IoT devices, exploring how these connections evolve over time and whether certificates are properly validated. Ren et al. [5] analysis the information exposure of devices in the United Kingdom and the United States. Inspired by their work, we introduce an analysis of Chinese devices and compare the result with other regions.
@@
    Data Collection Method Labels:SC (Self-Collection), CR (Crowdsourcing), SL (Device Simulation).
@@
IoT Dataset Construction. The IoT community has developed diverse datasets to facilitate researches in consumer IoT devices. Table 7 summarizes the most high-cited datasets. Among self-collected datasets, Mon(IoT)r [5] is the most cited and largest in device count, followed by UNSW  [8], YT  [18], Ours [6], and PingPong  [11]. Previous studies have provided valuable datasets for consumer IoT device research, but there are still some limitations. First, these datasets are collected before 2021, lacking recent consumer IoT traffic. Second, most existing datasets focus on device traffic from Europe and the United States, while traffic from Asia is rare. Third, current datasets do not fully consider the entire lifecycle of devices during the collection, restricting the analysis from the perspective of the lifecycle. In contrast, our dataset includes traffic with the entire lifecycle of the device, enabling researchers to conduct more fine-grained analyses of consumer IoT device traffic.
In summary, we construct the first Chinese consumer IoT traffic dataset that covers the entire lifecycle of a device, completing the research gap in the dataset for China. This dataset enables a more fine-grained analysis of traffic destinations and encryption practices and compares our measurements with those of other regions. By constructing the dataset and conducting measurements, this study provides new perspectives and insights for future research.
@@ 7 Conclusion
We collected 108 hours of traffic based on 77 consumer IoT devices across 12 categories and 38 brands. Through analysis, we draw the following key conclusions. First, 99.75% of traffic from Chinese consumer IoT devices is routed domestically, with only 0.25% directed overseas. In contrast, approximately 70% of the United Kingdom’s device traffic flowed overseas. This indicates minimal reliance on foreign services. However, a notable observation is that most devices depend on services from major corporations like Alibaba Cloud, Xiaomi, and Tencent, indicating that most devices rely on the services of these organizations, and these organizations may also obtain more user information from devices. Regarding encryption practices, Chinese consumer IoT devices demonstrate generally robust performance. The unencrypted traffic of Chinese devices peaked at 10%, lower than the other regions, and Chinese devices have less PII exposure in unencrypted traffic. Nevertheless, some devices still employ deprecated encryption protocols, insecure TLS certificate algorithms (e.g., SHA1, RSA-1024), and self-signed certificates. More critically, during device certificate replacement, certain devices exhibited an absence of certificate validation mechanisms, exposing communication to potential security and privacy risks.
We open-source our traffic collection guidance and make our dataset publicly available on the website [^2] https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to facilitate future research.
@@
This work was partly supported by the National Natural Science Foundation of China No.62102198 and the Key Program of the National Natural Science Foundation of China under No.62032012 and No.62432012.

### v3

@@ 5.5 Summary
This section measures the encryption practices of consumer IoT devices. The study reveals that Chinese manufacturers adopt safe encryption protocols for device communications, and encrypted traffic dominates most communications. However, some devices still use deprecated protocols, though their traffic constitutes a small proportion(7.6%), and some devices use proprietary protocols developed by the manufacturers rather than standardized protocols, bringing about certain security risks. Encouragingly, following firmware updates, 10 devices adopted more secure encryption protocols after the update, indicating a positive trend toward encryption protocol usage. Most devices use secure certificates, but prevalent self-signed certificates and insecure algorithms persist as potential risks. In terms of certificate verification, many devices have the problem of insecure certificate verification mechanisms, which may also bring potential risks. While direct PII exposure in unencrypted traffic is rare, inferred user preferences could enable targeted advertising. These findings indicate that the devices should adopt standard encryption protocols and reduce reliance on self-signed certificates to further minimize risks.
@@ 6 Comparisons on different regions
This section compares consumer IoT device traffic measurement results in China with those in other regions. We choose Ren et al.’s [9] and Paracha M. T. et al.’s [17] for comparison, as their studies align with our research purpose, focusing on traffic destinations and encryption analysis.
@@ 6.1 Traffic destinations
We have summarized the differences of traffic destinations between China and other regions in the table 7. In terms of traffic destinations, Chinese devices mainly communicate domestically when compared with consumer IoT devices in the United States and the United Kingdom.
@@
Other Regions 0.10% 0.10% 3.38% Greatbit Microsoft Microsoft
@@
party China (77) United States (46) United Kingdom(35) Tencent Kingsoft Kingsoft
@@
    The ratio of the United States and the United Kingdom are estimated in Ren et al.’s [9] work.
According to Ren et al. [9], the 35 devices in the United Kingdom communicated with 7 overseas regions. Approximately 70% of the United Kingdom’s device traffic flowed outside these regions (primarily to the United States). In contrast, although Chinese consumer IoT devices connected with 18 overseas countries, overseas traffic only accounted for 3% of total bytes. This highlights Chinese consumer IoT devices’ stronger reliance on domestic services than their United Kingdom counterparts. The devices in the United Kingdom partially rely on overseas cloud platforms and service providers (such as AWS, Google Cloud, and Akamai), resulting in frequent cross-border transmission of their data to overseas data centres. This increases the risks of data leakage. In contrast, most data transmission activities of Chinese consumer IoT devices occur within China, indicating a high degree of domestic service deployment. This result reflects the high degree of reliance of Chinese manufacturers on domestic services.
Regarding commonly contacted organizations, Chinese consumer IoT devices mainly interact with domestic providers. In Ren et al.’s study [9], the United Kingdom devices are mainly connected to the United States organizations, such as Amazon, Google, and Akamai. In our analysis, except for Google, Chinese devices primarily communicated with domestic organizations such as Alibaba Cloud, Xiaomi, and Tencent. These providers offer diverse services, including IoT platform services, cloud storage services, and video streaming and media content (See Section 4.2 for details).
@@ 6.2 Encryption analysis
Encryption practices vary regionally, but unencrypted traffic proportions remain low overall. Ren et al. [9] proposed entropy-based encryption detection, reporting unencrypted traffic proportions below 13% for devices in the United States and the United Kingdom. Using similar methods, we come to a similar conclusion: the percentage of unencrypted traffic of consumer IoT devices in China is low, with a peak of 11%.
Regarding encryption protocol usage, TLS 1.2 remains the most widely adopted, while TLS 1.3 adoption is still limited. Paracha M. T. et al. [17] analyzed 40 consumer IoT devices in the United States in 2021 and found that TLS 1.2 is the dominant protocol, with minimal use of TLS 1.3 and deprecated versions. Our findings are similar to these results. Among 77 Chinese devices, 81% utilized TLS 1.2, while 7.6% still relied on insecure protocols (TLS 1.1/SSL), and 14% employed proprietary protocols developed by manufacturers. For certificate validation, Paracha M. T. et al. [17] examined 32 devices and found that 11 were vulnerable to certificate replacement. In our study, 17 out of 40 devices generated error messages when the certificate was replaced, while 23 devices exhibited no errors or anomalies, indicating that certificate verification remains a security concern.
In terms of PII exposure, Ren et al.’s study [9] identified PII in unencrypted traffic, including MAC addresses, UUIDs, device IDs, geolocation data, and user-assigned device names. In contrast, Chinese devices exposed minimal PII, only singer names in the speaker’s traffic, indirectly revealing music preferences. This indicates stronger privacy protection in Chinese devices’ unencrypted traffic.
@@ 6.3 Summary
A clear regional difference is observed in the traffic destinations and encryption practices. Chinese consumer IoT devices are more reliant on domestic service providers, in contrast to devices in the United Kingdom, which mainly rely on international cloud platforms such as Amazon. This suggests that China has developed a more self-reliant and locally integrated IoT ecosystem, with limited dependence on international platforms. Encryption practices and certificate implementations are generally consistent across regions, ensuring relative security, but a certain proportion of devices still use deprecated protocols, which could pose potential security and privacy risks. Meanwhile, there are weaknesses with certificate verification in devices from different regions, posing a security concern. The PII contained in the unencrypted traffic of Chinese consumer IoT devices is less than that of their counterparts in the United Kingdom and the United States.
@@ 7 Related work
The measurement of IoT traffic. Researchers have conducted a lot of measurements on IoT traffic, mainly focusing on security and privacy. Security measurements involve mapping IoT backend servers [21], analyzing the impact of botnets on devices [30, 31], and investigating data leakage and device vulnerabilities through companion apps [33, 35, 36, 37]. Said et al. [21] mapped IoT backend servers by analyzing ISP passive traffic data, revealing relationships between backend providers. Arman et al. [30] assessed the impact of two ISP security policies on the Mirai botnet. Stephen et al. [31] studied the Hajime botnet’s effect on devices using active scanning and passive DNS backscatter traffic collection. These studies primarily focus on device-level evaluations while overlooking the user perspective. Privacy measurements include traffic destination identification [34], user data collection analysis [27, 28], and the effectiveness of existing privacy protection policies [29]. Mandalari et al. [34] proposed an automated method to identify IoT traffic destinations, distinguishing between essential and non-essential ones. Aniketh et al. [27] examined privacy leaks in local communication models using standard protocols. Dubois et al. [28] investigated how, when, and why smart speakers record external audio, while Anna Maria et al. [29] tested the effectiveness of existing privacy protection measures. Compared with these studies, our work directly analyzes IoT device security practices by evaluating traffic destinations and encryption practices. Ren et al. [9] and Paracha et al. [17] have similar research purposes to ours. Paracha et al. [17] evaluated the security of TLS connections in IoT devices, exploring how these connections evolve over time and whether certificates are properly validated. Ren et al. [9] analysis the information exposure of devices in the United Kingdom and the United States. Inspired by their work, we introduce an analysis of Chinese devices and compare the results with other regions.
@@
    Data collection method labels: SC (Self collection), CR (Crowd sourcing), SL (Device simulation).
@@
IoT dataset construction. The IoT community has developed diverse datasets to facilitate research in consumer IoT devices. Table 8 summarizes the most highly cited datasets. Among self-collected datasets, Mon(IoT)r [9] is the most cited and largest in device count, followed by UNSW  [12], YT  [22], Ours [10], and PingPong  [15]. Previous studies have provided valuable datasets for consumer IoT device research, but there are still some limitations. First, these datasets were collected before 2021, lacking recent consumer IoT traffic. Second, most existing datasets focus on device traffic from Europe and the United States, while traffic from Asia is rare. Third, current datasets do not fully consider the entire lifecycle of devices during the collection, restricting the analysis from the perspective of the lifecycle. In contrast, our dataset includes traffic with the entire lifecycle of the device, enabling researchers to conduct more fine-grained analysis of consumer IoT device traffic.
In summary, we construct the first Chinese consumer IoT traffic dataset that covers the entire lifecycle of devices, completing the research gap in the dataset for China. This dataset enables a more fine-grained analysis of traffic destinations and encryption practices, and compares our measurements with those of other regions. By constructing the dataset and conducting measurements, this study provides new perspectives and insights for future research.
@@ 8 Conclusion
We collected 108 hours of traffic based on 77 consumer IoT devices across 12 categories and 38 brands. Through analysis, we draw the following key conclusions. First, 99.75% of traffic from Chinese consumer IoT devices is routed domestically, with only 0.25% directed overseas. In contrast, approximately 70% of the United Kingdom’s device traffic flowed overseas. This indicates minimal reliance on overseas services. However, a notable observation is that most devices depend on services from major corporations like Alibaba Cloud, Xiaomi, and Tencent, indicating that most devices rely on the services of these organizations, and these organizations may also obtain more user information from devices. Regarding encryption practices, Chinese consumer IoT devices performed relatively well. The unencrypted traffic of Chinese devices peaked at 11%, lower than the other regions, and Chinese devices have less PII exposure in unencrypted traffic. Nevertheless, some devices still employ deprecated encryption protocols, insecure TLS certificate algorithms (e.g., SHA1, RSA-1024), and self-signed certificates. More critically, during device certificate replacement, certain devices exhibited an absence of certificate validation mechanisms, exposing communication to potential security and privacy risks.
We have open-sourced our traffic collection guidance and make our dataset publicly available on the website [^2] https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to facilitate future research.
@@
This work was partly supported by the National Natural Science Foundation of China No.62102198, the Fundamental Research Funds for the Central Universities No.079-63253221, and the Key Program of the National Natural Science Foundation of China under No.62032012 and No.62432012.